# Pitfalls of Rule- and Model-based Verifiers – A Case Study on Mathematical Reasoning

### v1

@@ 5.1 Classification-RL Performance Mismatch
We incorporate dedicated open-source verifiers explicitly fine-tuned for verification tasks, including: (1) xVerify 0.5B and 3B (Chen et al., 2025), fine-tuned on approximately 190K examples from multiple benchmarks; (2) general-verifier 1.5B(Ma et al., 2025), trained on a wide range of disciplines, including mathematics. (3) R1-Distill-Verifier-1.5B, a custom verifier we develop through rejection fine-tuning (Yuan et al., 2023) as detailed in Appendix G. The objective of this training is to reduce overthinking and encourage the model to generate more concise and focused outputs. It is worth noting that xVerify is a discriminative verifier, meaning it directly outputs a final judgment without intermediate reasoning. In contrast, other verifiers are generative, producing a chain-of-thought reasoning process before arriving at their final decision. For all trained verifiers, we apply an improved prompting strategy that includes the original question to provide additional context for verification. The static evaluation results for these verifiers are summarized in Table 1.
Static evaluation does not reflect long-term RL training.
As shown in Table 1, off-the-shelf verifiers trained on large-scale datasets significantly outperform general-purpose models. Among them, general-verifier achieves the highest performance, with a precision of 0.90 and a recall of 0.86. However, Table 2 reveals that the performance gap between the “general-verifier” model and “DS-R1-Distill-Qwen-1.5B” model is relatively small. This suggests that strong static evaluation metrics do not necessarily translate into superior performance during long-term RL training. Furthermore, our trained verifier, R1-Distill-Verifier-1.5B, also shows substantial gains over its base model, improving average recall from 0.49 to 0.62 and precision from 0.68 to 0.73 in static evaluation. Intuitively, we expect these improvements to translate into superior performance during dynamic RL training. However, we observe a counterintuitive phenomenon: as shown in the bottom right of Figure 1, after long-term RL training, the training reward experiences a significant surge at around 450 iterations. This anomaly leads us to investigate the occurrence of reward hacking, where the model optimizes for the reward signal in unintended ways that do not align with genuine performance improvements.
@@ 5.2 Verifier Under Siege: Reward Hacking in RL Training
To assess whether the rule-based or hybrid verifier provides an accurate reward signal and to detect potential reward hacking, we employ GPT-4o (Hurst et al., 2024) as an oracle during RL training. At each saved checkpoint, we randomly sample 1,000 queries from the training data, generate responses, and evaluate their correctness using GPT-4o. We then calculate the corresponding oracle reward. By analyzing the deviation between the training reward and the oracle reward, we gain valuable insights into both the effectiveness of the verifiers and the occurrence of reward hacking.
@@
Figure 1 (Right) plots the training reward against the oracle reward for different verifiers during RL training on DeepscaleR. Notably, after approximately 450 training iterations, the training reward using R1-Distill-Verifier-1.5B diverges significantly from the oracle reward provided by GPT-4o, while other methods maintain close alignment. This indicates that despite its strong static performance, R1-Distill-Verifier-1.5B becomes compromised during dynamic RL training, leading to a drop in evaluation accuracy and eventual training collapse, as shown in Figure 1 (Left). In contrast, the untrained verifier, R1-Distill-Verifier-1.5B, and the rule-based verifier do not exhibit such instability. These findings motivate our further investigation into verifier robustness in §6.
@@
Through analysis of the hacking phenomena, we observed that most of the attacks targeting R1-Distill-Verifier-1.5B fall into two patterns: Single Symbol and Gibberish. As shown in Figure 9 and Figure 10 in Appendix H, the policy model exploits vulnerabilities in the verifier during training by outputting either a single simple character (such as “{” ) or long sequences of meaningless text to bypass the verifier. Our observations are consistent with those of  Baker et al. (2025), indicating that although introducing a model-based verifier effectively increases the verifier’s flexibility, it implicitly raises the complexity of the environment and consequently reduces its robustness. Therefore, studying and improving the robustness of verifiers is of critical importance.
@@ 6 Probing Verifier Robustness with Hacking Patterns
Table 4: Success rates (%) of representative hacking patterns against verifiers. A lower success rate indicates that the model is less susceptible to hacking pattern attacks (i.e., lower is better). This table presents the success rates of selected representative hacking patterns, along with the overall average success rate. “DS” denotes DeepSeek, and for Qwen series models, the “instruct” suffix is omitted for clarity. Full results for all patterns are provided in Table 8 and Table 9.
@@
In §5, we find that static evaluation on the verification task fails to capture the verifier’s influence on long-term RL. Moreover, the trained verifier becomes increasingly vulnerable to hacking patterns over time. To better understand this vulnerability, we now perform a detailed robustness analysis of model-based verifiers. Building on the hacking patterns identified in §5.2, we design a broader set of attack patterns – ranging from simple gibberish inputs to more sophisticated adversarial prefixes. We then evaluate the effectiveness of these attacks across multiple model-based verifiers.
@@ 6.2 Analysis
Table 4 presents the success rates of various hacking patterns against different model-based verifiers, revealing that most verifiers are highly susceptible to these attacks. Strikingly, even simple patterns, such as inserting empty symbols (e.g., “{”) or injecting gibberish text, can effectively compromise many well-trained verifiers. Notably, our trained R1-Distill-Verifier-1.5B becomes more susceptible to such attacks after training. For instance, its susceptibility to adversarial prefixes rises from 21.7 (as seen in DeepSeek-R1-Distill-Qwen-1.5B) to 35, aligning with the observations discussed in §5.
@@ 7 Discussion
In this paper, we conduct a comprehensive analysis of rule-based and model-based verifiers within reinforcement learning for mathematical reasoning tasks. Our findings reveal critical pitfalls in both approaches: rule-based verifiers suffer from significant false negatives, particularly as policy models grow stronger, whereas model-based verifiers, despite higher accuracy in static evaluation, are notably vulnerable to reward hacking. This vulnerability results in inflated training rewards that fail to reflect genuine model performance, undermining the reliability of RL training outcomes. Future work should focus on developing robust verification systems that maintain accuracy without sacrificing robustness, thereby enhancing the reliability and effectiveness of reinforcement learning systems for complex reasoning tasks.
@@
One limitation of this paper is that our investigation primarily focuses on mathematical reasoning, a relatively well-defined domain. Although reinforcement learning with verifiable rewards has demonstrated utility across a broader range of complex tasks, including coding and agentic reasoning (Luo et al., 2025b; Liu and Zhang, 2025; Wang et al., 2025; Zheng et al., 2025), we think that in these broader domains, where RL is applied in much more complex environments, ensuring effectiveness and robustness of verifiers becomes even more critical. The insights and methods developed in our work may also be valuable for improving performance in these more challenging settings.
Appendix A Details of Verifier Evaluation Dataset Construction
In §3.1, we frame our dataset as a static classification task to assess the ability of verifiers to determine whether model responses align with a provided ground-truth answer. We use GPT-4o [Hurst et al., 2024] as an annotator to generate ground-truth labels, evaluating each response against the target answer according to the prompt shown in Figure 4.
@@ Appendix A Details of Verifier Evaluation Dataset Construction
Appendix B Detailed Results of Rule-based Verifiers Across Datasets
@@
Appendix C Detailed Evaluation Setting for Model-based Verifiers
@@
For untrained verifiers, including (1) Short-CoT models: Qwen-2.5-instruct (1.5B and 7B) [Yang et al., 2024b] and Qwen-2.5-math-instruct (1.5B and 7B) [Yang et al., 2024a]. (2) R1-style long-CoT models: DeepSeek-R1-Distill-Qwen (1.5B and 7B) [DeepSeek-AI et al., 2025], we employed a simplified prompt format during evaluation, providing only the ground truth and the model-generated answer to reduce overthinking. For the trained verifier, we apply an improved prompting strategy that includes the original question to provide additional context for verification. Prompts that include and exclude the original question for these verifiers are detailed in Figure 6.
@@
Most verifiers used greedy decoding during evaluation. An exception was made for the R1-style Long-CoT models (including our trained R1-Distill-Verifier-1.5B), for which we followed the settings of DeepSeek-AI et al. [2025], applying temperature = 0.6 and top-p = 0.95 to reduce output repetition.
Appendix D Detailed Results of Model-based Verifiers and Hybrid Verifiers
@@
Appendix E Training And Evaluation Details of Reinforcement Learning
@@
We use Verl [Sheng et al., 2024] as the RL training framework and implement the model-based verifier within the HybridEngine architecture. HybridEngine efficiently partitions models and dynamically switches between training and inference modes, significantly improving GPU utilization and reducing communication overhead during RL training. Building on this capability, we extend HybridEngine to the model-based verifier, allowing it to be offloaded from GPUs during idle periods. For alternative implementations – such as assigning the verifier to dedicated GPUs or deploying it as a standalone server [Su et al., 2025, Ma et al., 2025] – we minimize contention between the policy model and the model-based verifier, further enhancing GPU efficiency.
@@
We train our models using the Verl framework [Sheng et al., 2024]. The Training uses a prompt batch size of 1,024, generating 8 rollouts per prompt with a maximum rollout length of 8,192 tokens. We apply a mini-batch size of 256 for updates. The sampling temperature is set to 1.0 by default. Following Yu et al. [2025], we set the clip_high ratio to 0.28, maintain clip_low at 0.2, and set the KL coefficient to 0. We used the same training prompt as  Zeng et al. [2025a].
@@
We build our evaluation script based on  Yang et al. [2024a], using temperature of 1.0 and topp of 0.7 and a maximum generation length of 16K tokens. To ensure consistency, we adopt the same prompt template used during training. For most benchmarks, we report pass@1 results. However, for AIME 2024, which contains fewer problems, we report both pass@1 and average accuracy (avg@32), computed over 32 generated samples per problem.
@@
Appendix F Details Results of Reinforcement Learning
Training Dynamic using general-verifier
@@ Appendix F Details Results of Reinforcement Learning
Appendix G Training Details for R1-Distill-Verifier-1.5B
To reduce overthinking and encourage more concise, focused outputs, we fine-tune DeepSeek-R1-Distill-Qwen-1.5B using rejection fine-tuning [Yuan et al., 2023]. Specifically, we sample 1K queries from the DeepscaleR dataset (non-overlapping with the evaluation set described in §3.1). For each query, we generate eight candidate responses from DeepSeek-R1-Distill-Qwen-32B and use GPT-4o [Hurst et al., 2024] as an annotator to assess whether each response aligns with the ground-truth answer. We then sample eight candidate responses from DeepSeek-R1-Distill-Qwen-1.5B. Responses that do not match GPT-4o’s judgment or are duplicates are filtered out, yielding approximately 20K examples for fine-tuning. The model is fully fine-tuned using a learning rate of 1e-4 for 3 epochs.
@@
Figure 8: The training and evaluation curves of RL using general-verifier, with the x-axis representing training iterations in all plots. Left illustrates the evaluation accuracy averaged over multiple benchmarks, including GSM8K, MATH500, Minerva Math, OlympiadBench, AIME24, and AMC23. Right depicts changes in reward values during training. The “training rewards” indicate the rewards provided by the corresponding reward system to the policy model, whereas the “oracle rewards” represent rewards the model receives when judged by combining with GPT-4o. We provide a detailed breakdown of evaluation results in Table 2.
Appendix H Analysis of Hacking Pattern During RL Training

### v2

@@ 5.1 Classification-RL Performance Mismatch
We incorporate dedicated open-source verifiers explicitly fine-tuned for verification tasks, including: (1) xVerify 0.5B and 3B (Chen et al., 2025), fine-tuned on 190K examples from multiple benchmarks; (2) general-verifier 1.5B(Ma et al., 2025), trained on diverse disciplines, including mathematics. (3) R1-Distill-Verifier-1.5B, a custom verifier we develop through rejection fine-tuning (Yuan et al., 2023) as detailed in Appendix K. The objective of this training is to reduce overthinking and encourage the model to generate more concise and focused outputs. It is worth noting that xVerify is a discriminative verifier that outputs direct judgments, while the others are generative, producing chain-of-thought reasoning. For all trained verifiers, we apply an improved prompting strategy that includes the original question to provide additional context for verification. The static evaluation results for these verifiers are summarized in Table 1.
Static evaluation does not necessarily reflect long-term RL training.
As shown in Table 1, the verifiers trained on labeled classification data significantly outperform general-purpose models. Our trained verifier, R1-Distill-Verifier-1.5B, shows substantial gains over its base model, improving average recall from 0.49 to 0.62 and precision from 0.68 to 0.73 in static evaluation. Intuitively, we expect these improvements to translate into superior performance during dynamic RL training. However, we observe a counterintuitive phenomenon: as shown in the bottom right of Figure 1, after long-term RL training, the training reward surges at around 450 iterations. Despite this increase, the best evaluation results (Table 2) show little improvement over the rule-based verifier (55.6 vs. 55.0). Moreover, experiments on the Skywork-OR1 dataset reveal an even clearer degradation, with performance dropping from 58.7 to 55.5 when using our trained verifier, as shown in Figure 9 in Appendix I. These anomalies point to the presence of reward hacking, where the model exploits weaknesses in the reward signal to inflate rewards without genuine performance improvements.
@@ 5.2 Verifier Under Siege: Reward Hacking in RL Training
To assess whether the rule-based or hybrid verifier provides an accurate reward signal and to detect potential reward hacking, we employ GPT-4o (Hurst et al., 2024) as an oracle during RL training. At each checkpoint, we sample 1,000 training queries, generate responses, and assess correctness with GPT-4o to compute the oracle reward. By analyzing the deviation between the training reward and the oracle reward, we gain valuable insights into both the effectiveness of the verifiers and the occurrence of reward hacking.
@@
Figure 1 (Right) plots the training reward against the oracle reward for different verifiers during RL training on DeepscaleR. Notably, after approximately 450 training iterations, the training reward using R1-Distill-Verifier-1.5B diverges significantly from the oracle reward provided by GPT-4o, while other methods maintain close alignment. The oracle reward further reveals a steep decline toward the end of training. This indicates that despite its strong static performance, R1-Distill-Verifier-1.5B becomes compromised during dynamic RL training, leading to a drop in evaluation accuracy and eventual training collapse, as shown in Figure 1 (Left). In contrast, the untrained verifier, R1-Distill-Verifier-1.5B, and the rule-based verifier do not exhibit such instability. These findings motivate our further investigation into verifier robustness in §6.
@@
Most exploits against R1-Distill-Verifier-1.5B fall into two patterns: Single Symbol and Gibberish. As shown in Figure 11 and Figure 12 in Appendix L, the policy model exploits vulnerabilities in the verifier by outputting either a single simple character (such as “{” ) or long sequences of meaningless text to bypass the verifier. Consistent with  Baker et al. (2025), these results suggest that although introducing a model-based verifier effectively increases the verifier’s flexibility, it implicitly raises the complexity of the environment and reduces its robustness. Therefore, studying and improving the robustness of verifiers is of critical importance.
Reward Hacking Beyond Math.
To test whether these vulnerabilities generalize, we further conducted RL experiments on Skywork-OR1 (He et al., 2025) (math domain) and WebInstruct-Verified (Ma et al., 2025) (general science). As detailed in Appendix I and Appendix J, reward hacking persists across both domains: trained model-based verifiers remain susceptible, underscoring that the challenge is not confined to a single dataset but inherent to broader reasoning tasks.
Notably, while this section focuses on reward hacking in fine-tuned verifiers, one might assume that general LLM verifiers are relatively robust due to the RL improvements described in §4.3. However, in the next section, we show that even simple patterns can severely undermine both general and fine-tuned verifiers, revealing significant risks associated with relying on model-based verifiers.
@@ 6 Probing Verifier Robustness with Hacking Patterns
Table 3: Success rates (%) of representative hacking patterns against verifiers. A lower success rate indicates that the model is less susceptible to hacking pattern attacks (i.e., lower is better). This table presents the success rates of selected representative hacking patterns, along with the overall average success rate. “DS” denotes DeepSeek, and for Qwen series models, the “instruct” suffix is omitted for clarity. Full results for all patterns are provided in Table 9 and Table 10 in Appendix M.
@@
Motivated by our findings in §5, where the trained verifier exhibits increasing vulnerability to hacking patterns over time, we conduct a systematic probing study to expose risks faced by both untrained and fine-tuned verifiers. We argue that the evaluation of model-based verifiers should not only emphasize accuracy, but also robustness against adversarial manipulation. Building on the hacking patterns identified in §5.2, we construct a broader suite of attack strategies – ranging from simple gibberish inputs to more sophisticated adversarial prefixes. We then evaluate the effectiveness of these attacks across multiple model-based verifiers, enabling a more comprehensive assessment of their robustness under adversarial conditions.
@@ 6.2 Analysis
Table 3 reports the success rates of different hacking patterns across various model-based verifiers, showing that all generative verifiers – regardless of whether they are fine-tuned or not – are highly vulnerable to these attacks. Remarkably, even trivial manipulations, such as inserting empty symbols (e.g., “{”) or appending gibberish text, can reliably compromise most verifiers. Furthermore, our trained R1-Distill-Verifier-1.5B becomes even more fragile after training: its susceptibility to adversarial prefixes increases from 21.7 (observed in DeepSeek-R1-Distill-Qwen-1.5B) to 35, consistent with the trends identified in §5.
@@
Reward hacking during RL training.
As shown in Figure 1 (Right), we observe clear reward hacking when R1-Distill-Verifier-1.5B is used as the RL verifier, consistent with its vulnerability to simple attacks such as empty symbols (Table 3). Interestingly, DS-R1-Distill-Qwen-1.5B does not show reward hacking in RL experiments, yet Table 3 still reports abnormally high attack success rates. We hypothesize that this is because the policy models in our RL training are not strong enough to find and exploit these vulnerabilities of DS-R1-Distill-Qwen-1.5B. Importantly, we stress that base models are not inherently safe: even the simplest empty symbols attack can hack them at scale. This highlights the urgency of deeper investigations into verifier robustness, particularly in RL training with stronger models.
@@ 7 Discussion
In this paper, we conduct a comprehensive analysis of rule-based and model-based verifiers within reinforcement learning for mathematical reasoning tasks. Our findings reveal critical challenges in both approaches: rule-based verifiers suffer from significant false negatives, particularly as policy models grow stronger, whereas model-based verifiers, despite higher accuracy in static evaluation, are notably vulnerable to reward hacking. This vulnerability results in inflated training rewards that fail to reflect genuine model performance, undermining the reliability of RL training outcomes. Future work should focus on developing robust verification systems that maintain accuracy without sacrificing robustness, thereby enhancing the reliability and effectiveness of reinforcement learning systems for complex reasoning tasks.
@@
This paper primarily analyzes rule-based and model-based verifiers, highlighting their limitations and vulnerabilities. We view this as an important first step toward addressing the broader challenge of building trustworthy verifiers, and we hope future work will further advance this direction.
Appendix A The Use of Large Language Models
In this paper, large language models (LLMs) are used exclusively for language polishing. The entire research process, including ideation and all subsequent stages, was carried out without any assistance from LLMs.
Appendix B Details of Verifier Evaluation Dataset Construction
Prompt Format.
In §3.1, we frame our dataset as a static classification task to assess the ability of verifiers to determine whether model responses align with a provided ground-truth answer. We use GPT-4o (Hurst et al., 2024) as an annotator to generate ground-truth labels, evaluating each response against the target answer according to the prompt shown in Figure 4.
Justification of GPT-4o annotation.
@@ Appendix B Details of Verifier Evaluation Dataset Construction
Appendix C Technical Details about Rule-based Verifier
Below is a brief summary of the key differences between the rule-based verifier implementations mentioned in our work:
  • Verl Math Verifier: Implemented in the official VERL repository, this verifier is relatively simple and primarily based on string matching. It does not perform LaTeX compilation and therefore lacks the ability to evaluate mathematical equivalence at a semantic level.
  • Qwen-Math Verifier: Originally developed by the Qwen team as part of their math evaluation framework, and later adopted by the SimpleRL (Zeng et al., 2025a; b) team as an RL-compatible verifier. It supports LaTeX compilation, allowing it to handle higher-level mathematical equivalences more robustly.
  • HuggingFace Math Verifier: Introduced after Qwen-Math by the HuggingFace team, this verifier also incorporates LaTeX compilation, albeit with some differences in implementation details. In practice, its performance is generally considered to be on par with the Qwen-Math verifier.
Appendix D Detailed Results of Rule-based Verifiers Across Datasets
@@
Appendix E Detailed Evaluation Setting for Model-based Verifiers
@@
For untrained verifiers, including (1) Short-CoT models: Qwen-2.5-instruct (1.5B and 7B) (Yang et al., 2024a) and Qwen-2.5-math-instruct (1.5B and 7B) (Yang et al., 2024b). (2) R1-style long-CoT models: DeepSeek-R1-Distill-Qwen (1.5B and 7B) (DeepSeek-AI et al., 2025), we employed a simplified prompt format during evaluation, providing only the ground truth and the model-generated answer to reduce overthinking. For the trained verifier, we apply an improved prompting strategy that includes the original question to provide additional context for verification. Prompts that include and exclude the original question for these verifiers are detailed in Figure 6.
@@
Most verifiers used greedy decoding during evaluation. An exception was made for the R1-style Long-CoT models (including our trained R1-Distill-Verifier-1.5B), for which we followed the settings of DeepSeek-AI et al. (2025), applying temperature = 0.6 and top-p = 0.95 to reduce output repetition.
Appendix F Detailed Results of Model-based Verifiers and Hybrid Verifiers
@@
Appendix G Training And Evaluation Details of Reinforcement Learning
@@
We use Verl (Sheng et al., 2024) as the RL training framework and implement the model-based verifier within the HybridEngine architecture. HybridEngine efficiently partitions models and dynamically switches between training and inference modes, significantly improving GPU utilization and reducing communication overhead during RL training. Building on this capability, we extend HybridEngine to the model-based verifier, allowing it to be offloaded from GPUs during idle periods. For alternative implementations – such as assigning the verifier to dedicated GPUs or deploying it as a standalone server (Su et al., 2025; Ma et al., 2025) – we minimize contention between the policy model and the model-based verifier, further enhancing GPU efficiency.
@@
We train our models using the Verl framework (Sheng et al., 2024). The Training uses a prompt batch size of 1,024, generating 8 rollouts per prompt with a maximum rollout length of 8,192 tokens. We apply a mini-batch size of 256 for updates. The sampling temperature is set to 1.0 by default. Following Yu et al. (2025), we set the clip_high ratio to 0.28, maintain clip_low at 0.2, and set the KL coefficient to 0. We used the same training prompt as  Zeng et al. (2025a).
@@
We build our evaluation script based on  Yang et al. (2024b), using temperature of 1.0 and topp of 0.7 and a maximum generation length of 16K tokens. To ensure consistency, we adopt the same prompt template used during training. For Figure 1, all benchmarks are evaluated with a single random sampling. For AIME 2024 and AMC 2023, we additionally report stable results by averaging over 32 random samplings (Avg@32) in Table 2.
@@
Appendix H RL Experiments on DeepscaleR Dataset
@@
Figure 8: The training and evaluation curves of RL using general-verifier on DeepScaleR dataset, with the x-axis representing training iterations in all plots. Left illustrates the evaluation accuracy averaged over multiple benchmarks, including GSM8K, MATH500, Minerva Math, OlympiadBench, AIME24, and AMC23. Right depicts changes in reward values during training. The “training rewards” indicate the rewards provided by the corresponding reward system to the policy model, whereas the “oracle rewards” represent rewards the model receives when judged by combining with GPT-4o. We provide a detailed breakdown of evaluation results in Table 2.
Appendix I RL Experiments on Skywork-OR1 Dataset
We further perform detailed RL experiments on Skywork-OR1, following the basic experiment settings of the DeepscaleR dataset. We train Qwen-2.5-7B for 700 steps. For the verifiers, we include rule-based verifier, R1-Verifier-1.5B (hybrid) and general-verifier (hybrid), respectively. The detailed results are shown in Figure 9 and Table 6. At step 400, the training reward using R1-Distill-Verifier-1.5B diverges significantly from the oracle reward provided by GPT-4o, indicating the presence of reward hacking. Overall, in Table 6, the results show that the general-verifier achieves higher RL performance than the rule-based verifier, primarily due to its higher recall rate. However, with the R1-Distill-Verifier-1.5B verifier, where reward hacking appears, the performance is significantly lower than other settings, further validating that our finding can generalize to other mathematical datasets.
Figure 9: Changes in reward values during training on Skywork-OR1. The “training rewards” indicate the rewards provided by the corresponding reward system to the policy model, whereas the “oracle rewards” represent rewards the model receives when judged by combining with GPT-4o.
Table 6: Detailed performance of models trained on Skywork-OR1 across multiple benchmarks. The best result from each run is reported. Blue lines indicate models trained with a hybrid verifier without evidence of reward hacking, while pink lines indicate runs where reward hacking is detected. “HF” represents HuggingFace Math Verifier. Training curves for these models are presented in Figure 9.
Model GSM8K
MATH
500
Minerva
Math
Olympiad
Bench
AIME24
(Avg@32)
AMC23
(Avg@32)
Avg.
Qwen-2.5-7B-SimpleRL-Zoo 91.7 78.2 38.6 40.4 15.6 54.9 53.2
Qwen-2.5-7B 88.2 64.6 25.7 30.1 0.3 36.9 41.0
      $\hookrightarrow$ + Skywork-OR1 & HF verifier 93.2 82.6 38.6 48.9 22.6 66.1 58.7
          $\hookrightarrow$ + general-verifier 93.3 86.2 35.3 52.9 23.4 72.7 60.6
          $\hookrightarrow$ + R1-Distill-Verifier-1.5B verifier 93.0 79.6 34.9 45.6 16.8 63.3 55.5
Appendix J Experiments on WebInstruct-Verified Dataset
To assess the generality of our findings beyond the math domain, we conduct additional static evaluation and RL experiments on WebInstruct-Verified (He et al., 2025), which spans broader domains such as physics and finance. We present the static evaluation results and RL outcomes below.
Static Evaluation.
We first evaluate the static performance of rule-based verifiers using DeepSeek-R1-Distill-Qwen-32B as the policy model, following the setup in Figure 3. The results are summarized in Table 7. Here, rule-based verifiers perform even worse, with recall dropping below 0.6, highlighting their limited adaptability to diverse and less structured answer formats.
Table 7: Performance of different rule-based verifiers on WebInstruct-Verified dataset. Evaluations are conducted on a subset of the static evaluation dataset sampled from Deepseek-R1-Distill-Qwen-32B.
Verifier Precision Recall
VERL Verifier 1.00 0.35
Qwen-Math Verifier 1.00 0.58
HuggingFace Verifier 0.98 0.47
RL Experiments.
We further train Qwen-2.5-7B on WebInstruct-Verified using a hybrid setup with R1-Verifier-1.5B. Performance is evaluated on GPQA-Diamond (Rein et al., 2024). As shown in Figure 10 (Right), the training reward (from the verifier) diverges from the oracle reward (from GPT-4o) after around 200 steps, with the gap reaching approximately 0.2. Beyond this point, as illustrated in Figure 10 (Left), downstream performance on GPQA-Diamond stagnates and even declines, confirming the presence of reward hacking. These findings mirror the patterns observed in the math domain, reinforcing the broader validity of our conclusions.
Figure 10: Left illustrates the evaluation accuracy over GPQA-Diamond. Right depicts changes in reward values during training. The “training rewards” indicate the rewards provided by the corresponding reward system to the policy model, whereas the “oracle rewards” represent rewards the model receives when judged by combining with GPT-4o.
Appendix K Training Details for R1-Distill-Verifier-1.5B
To reduce overthinking and encourage more concise, focused outputs, we fine-tune DeepSeek-R1-Distill-Qwen-1.5B using rejection fine-tuning (Yuan et al., 2023). Specifically, we sample 1K queries from the DeepscaleR dataset (non-overlapping with the evaluation set described in §3.1). For each query, we generate eight candidate responses from DeepSeek-R1-Distill-Qwen-32B and use GPT-4o (Hurst et al., 2024) as an annotator to assess whether each response aligns with the ground-truth answer. We then sample eight candidate responses from DeepSeek-R1-Distill-Qwen-1.5B. Responses that do not match GPT-4o’s judgment or are duplicates are filtered out, yielding approximately 20K examples for fine-tuning. The model is fully fine-tuned using a learning rate of 1e-4 for 3 epochs.
Appendix L Analysis of Hacking Pattern During RL Training